Budgeting Cash Collections
Spencer Consulting, which invoices its clients on terms 2/10, n/30, had credit sales for May and June of $112,000 and $128,000, respectively. Analysis of Spencer's operations indicates that the pattern of customers' payments on account is as follows (percentages are of total monthly credit sales):

Receiving Discount Beyond Discount Period Totals
In month of sale 50% 20% 70%
In month of following sale 15% 10% 25%
Uncollectible accounts, returns, and allowances 5%
100%

Determine the estimated cash collected on customers' accounts in June.

$Answer
Answer #1

SOLUTION:

1) 2/10 means if you pay the amount in 10 days you will get 2% discount.

Estimated Cash Collected on Customer Account for the month of June

Working Notes:

In the given question it is given that sales for the month of May is $1,12,000

Therefore the Payment for the sales will be received in the following Manner

in month of may :

1) Receiving Discount of 2%, 50% of the sales amount will be collected, Hence amount = 1,12,000*50%= 56000 but the amount receivable is 56000-2%= 54880 (cash received)

2) And after 10 Days for the May month sales you will receive 20% Payment, Hence amount = (1,12,000-56000)*20%=11200

Therefore May month Collection is 54880 + 11200. =66,080/-

Balance amount which is still receivable from customer for the month of May is 1,12,000-56000-11200= 44800 (Note-1)

Note 1) Here 56000 is supposed to be taken that is before discount while calculating amount payable by customer and not what customer has actually paid to us i.e. 54880.

Amount receivable for the month of June:

Two Amount will be receivable i.e. May Sales balance amount and June sales

a) May Amount- 1) 15% of the amount is received with discount, so amount payable by customer is 44800*15%=6720 on which 2% discount will be given.

Therefore net amount is 6720-2%= 6585.6

2) 10% of the amount is received: i.e. (44800-6720)*10%=3808.

Therefore total amount received in June for May sales is = 6585.6+3808=10393.6

b) June amount- 1) Sales of June is Rs 128000 out of which- 50% of the amount will be received after providing 2% discount i.e. (128000*50%)-2%= 62720

2) From sales of June after 10 days period of availing discount you will receive 20% of the amount from the customer which comes to = (128000-64000)*20%=12800

Therefore from sales of June amount received is 62720+12800=75,520/-

Therefore total collection of June is 75550+10393.6=85913.6/- from the customer.